# Abdulrahman Alswaidan and Jeffrey D. Varner R.F. Smith School of Chemical and Biomolecular Engineering

### 06 Mar 2026 v1

@@ 5 Experiments
The synthetic experiments above validated the sampler’s statistical properties on controlled data. We tested whether it could generate structured, recognizable outputs from real-world patterns and compared it against six baselines. We selected $K=100$ images of the digit “3” from the MNIST training set [18], flattened each to a vector in $\mathbb{R}^{784}$, and normalized to unit norm to form the memory matrix $\mathbf{X}\in\mathbb{R}^{784\times 100}$ (load ratio $K/d\approx 0.13$). Because the energy landscape at high $\beta$ is deeply multimodal, a single long chain can become trapped in one basin for the entire run. We therefore launched 30 independent chains, each initialized near a different randomly chosen stored pattern with a small Gaussian perturbation ($\sigma_{\mathrm{init}}=0.01$). Each chain ran Algorithm 1 at $\beta=2000$ with step size $\alpha=0.01$ for T=5, [^0] T=5{,}000 iterations; after discarding the first 2,0002{,}000 steps as burn-in, thinning every 100th iteration to obtain 30 post-burn-in snapshots, and sub-sampling 5 evenly spaced snapshots per chain, we retained $30\times 5=150$ generated images that collectively span 30 distinct energy basins.
The choice of $\beta=2000$ follows from the per-step signal-to-noise ratio,
@@
where numerator and denominator are the gradient displacement and expected noise magnitude. The synthetic experiment placed the transition near $\mathrm{SNR}\approx 0.025$ ($\beta\approx 5$$10$, d=64d{=}64); inverting at d=784d{=}784 gives $\beta\approx 100$, and we set $\beta=2000$ ($\mathrm{SNR}=0.113$, $4\times$ the transition) to operate well inside the structured regime.
@@
Table 1: Quantitative comparison on MNIST digit “3” ($K=100$, $d=784$). $\mathcal{N}$ and $\bar{\mathcal{D}}$ are higher-is-better ($\uparrow$); $\bar{E}$ is lower-is-better ($\downarrow$). Values are mean $\pm$ SE across 30 chains. Two SA rows demonstrate the temperature knob: $\beta{=}2000$ ($\mathrm{SNR}{=}0.113$, structured retrieval) and $\beta{=}200$ ($\mathrm{SNR}{=}0.036$, generation). Energy is positive at $\beta{=}200$ because samples explore off the attractor manifold.
@@
Table 1 and Figure 2 summarized the results. Each non-Langevin baseline failed on at least one axis: bootstrap by construction ($\mathcal{N}{=}0$); Gaussian perturbation by negligible noise ($\sigma{\approx}0.003$, indistinguishable from bootstrap); and random convex combination by mode collapse ($\bar{\mathcal{D}}{=}0.008$). Among static baselines, the VAE (latent dim 8, trained on the same K=100K{=}100 patterns) is the strongest: $\mathcal{N}{=}0.214\pm 0.005$, $\bar{\mathcal{D}}{=}0.441\pm 0.008$, surpassing GMM-PCA on both axes. At $\beta{=}200$ (SNR==0.036, the generation regime), SA dominates every baseline, including the best learned model (VAE), on both generation metrics simultaneously: novelty $0.548\pm 0.002$ ($2.6{\times}$ VAE, $2.8{\times}$ GMM-PCA), diversity $0.885\pm 0.002$ ($2.0{\times}$ VAE, $2.1{\times}$ GMM-PCA); energy rises above zero as samples explore off the attractor manifold (see ). MALA’s 99.2% acceptance rate at both $\beta$ values confirms negligible ULA bias across regimes. The SNR transition band (0.020.020.030.03) thus delimits two qualitatively distinct regimes in the same algorithm: structured retrieval ($\beta{=}2000$, crisp digits near stored patterns) and genuine generation ($\beta{=}200$, blurry-but-recognizable novel digits). To verify that the high diversity at $\beta{=}200$ reflects genuine basin-crossing rather than multi-chain initialization, we ran a single long chain (T=50, [^0] T{=}50{,}000) from a fixed seed; single-chain diversity reached $0.796$, exceeding the 30-chain $\beta{=}2000$ value of $0.600$ (Table 5, Appendix F).
@@
Appendix E confirmed the same ranking on digits “1” and “8” (SA: $\mathcal{N}{\approx}0.152$$0.153$, $\bar{\mathcal{D}}{\approx}0.557$$0.591$; VAE: $\mathcal{N}{\approx}0.180$$0.209$, $\bar{\mathcal{D}}{\approx}0.397$$0.408$; MALA acceptance 99.2%): the VAE remains the strongest non-Langevin baseline on novelty but is far below both Langevin methods on diversity across all digit classes. Appendix G used S&P 500 log-returns (d=424d{=}424, K=2,766K{=}2{,}766) to characterize the novelty–fidelity trade-off on real data: SA achieved novelty $0.768\pm 0.001$ (bootstrap: $0.000$) while bootstrap led on marginal fidelity ($99.3\%$ vs. $64.2\%$ KS pass); volatility clustering was not reproduced, a principled consequence of fixed-$\beta$ equilibrium sampling. Appendix H scales to face images at d=4,096d{=}4{,}096: the SNR rule prescribes $\beta{=}10{,}000$; SA achieves novelty $0.159\pm 0.000$ and diversity $0.293\pm 0.001$, with ranking identical to MNIST across a $5.2{\times}$ dimension increase.
@@ 6 Discussion and Conclusion
The experiments validated the core claim: Langevin dynamics on the modern Hopfield energy converts deterministic attention into a principled stochastic sampler governed by a single temperature parameter. The temperature spectrum (Fig. 1a) exhibited a smooth phase transition between retrieval and generation, the continuous analogue of the classical Hopfield/Boltzmann duality [1] lifted to continuous states and Langevin dynamics. The convergence experiment (Fig. 1b) confirmed that Algorithm 1 reached the correct Boltzmann target, consistent with ULA theory [7]. The load-ratio sweep (Fig. 3) showed that generation quality degraded gracefully as $K/d$ increased, providing empirical evidence linking modern Hopfield capacity to generation quality. The MNIST experiment (Table 1, Fig. 2) demonstrated that these properties carried over to real images: the sampler generated novel, low-energy digit images that no static baseline could approach, including a VAE trained on the same K=100K{=}100 patterns ($\mathcal{N}{=}0.214$, $\bar{\mathcal{D}}{=}0.441$, the strongest non-Langevin result): at $\beta{=}200$, SA led the VAE by $2.6{\times}$ on novelty and $2.0{\times}$ on diversity. SA matched the Metropolis-corrected MALA baseline, confirming ULA’s discretization bias was negligible at $\alpha{=}0.01$. Appendices G and H extended these results to S&P 500 log-returns (d=424d{=}424) and face images (d=4,096d{=}4{,}096), confirming the SNR rule and the novelty–fidelity trade-off across a $5.2{\times}$ dimension range.
In summary, stochastic attention reuses the same primitive operations as a standard attention head and requires no learned score network, no training loop, and no contrastive objective. The analytic structure of the Hopfield energy (smoothness, dissipativity, and an exact closed-form gradient) provides convergence guarantees that generic energy-based samplers cannot offer without additional assumptions. Across all four experimental domains ($d\in\{64,\,784,\,424,\,4{,}096\}$), the transition from structured retrieval to diffuse generation consistently occurs near $\mathrm{SNR}=\sqrt{\alpha\beta/2d}\approx 0.02$$0.03$, giving a dimension-independent rule for targeting $\beta$: set $\beta=\mathrm{SNR}_{\mathrm{target}}^{2}\cdot 2d/\alpha$.
@@
The ULA discretization introduces bias scaling with $\alpha$; our MALA comparison (Table 1) shows a 99.2% acceptance rate at $\alpha{=}0.01$, confirming negligible bias at this step size. At larger $\alpha$, MALA (Algorithm 2, Appendix C) is the preferred variant. Mixing time grows with both $\beta$ and $K/d$, and quantitative scaling laws remain to be established. Each step costs $\mathcal{O}(NK)$ (two matrix–vector products plus a softmax), identical to a single attention head; for N=784N{=}784, K=100K{=}100 the full 30-chain experiment completes in under 10 minutes on a laptop CPU. The method also requires $\mathbf{X}$ to be known and fixed at sampling time; streaming or latent memory settings would require extension. We consider only single-head attention over a flat memory; extension to multi-head attention, hierarchical memories, or full encoder-decoder transformers is future work.

### 10 Mar 2026 v2

@@ 5 Experiments
The synthetic experiments above validated the sampler’s statistical properties on controlled data. We tested whether it could generate structured, recognizable outputs from real-world patterns and compared it against seven baselines. We selected $K=100$ images of the digit “3” from the MNIST training set [21], flattened each to a vector in $\mathbb{R}^{784}$, and normalized to unit norm to form the memory matrix $\mathbf{X}\in\mathbb{R}^{784\times 100}$ (load ratio $K/d\approx 0.13$). Because the energy landscape at high $\beta$ is deeply multimodal, a single long chain can become trapped in one basin for the entire run. We therefore launched 30 independent chains, each initialized near a different randomly chosen stored pattern with a small Gaussian perturbation ($\sigma_{\mathrm{init}}=0.01$). Each chain ran Algorithm 1 at $\beta=2000$ with step size $\alpha=0.01$ for T=5, [^0] T=5{,}000 iterations; after discarding the first 2,0002{,}000 steps as burn-in, thinning every 100th iteration to obtain 30 post-burn-in snapshots, and sub-sampling 5 evenly spaced snapshots per chain, we retained $30\times 5=150$ generated images that collectively span 30 distinct energy basins.
The choice of $\beta=2000$ follows from the phase transition analysis of Section 4.3. The per-step signal-to-noise ratio,
@@
where numerator and denominator are the gradient displacement and expected noise magnitude, undergoes a critical transition at $\mathrm{SNR}^{*}=\sqrt{\alpha/(2\sqrt{d})}$ (Eq. 14), the entropy inflection point of Proposition 3. At d=64d{=}64 and $\alpha{=}0.01$, this gives $\mathrm{SNR}^{*}=0.025$ ($\beta^{*}\approx 8$), consistent with the observed transition at $\beta\approx 5$$10$. For structured data the effective dimensionality is lower and the transition $\beta$ shifts accordingly; we set $\beta=2000$ ($\mathrm{SNR}=0.113$) to operate well inside the structured-retrieval regime.
@@
Table 1: Quantitative comparison on MNIST digit “3” ($K=100$, $d=784$). $\mathcal{N}$ and $\bar{\mathcal{D}}$ are higher-is-better ($\uparrow$); $\bar{E}$ is lower-is-better ($\downarrow$). Values are mean $\pm$ SE across 30 chains. Two SA rows demonstrate the temperature knob: $\beta{=}2000$ ($\mathrm{SNR}{=}0.113$, structured retrieval) and $\beta{=}200$ ($\mathrm{SNR}{=}0.036$, near transition edge; generation regime). Energy is positive at $\beta{=}200$ because samples explore off the attractor manifold. DDPM generates unstructured noise (max-$\cos=0.062$, indistinguishable from isotropic Gaussian); the model cannot learn the data distribution from K=100K{=}100 training images.
@@
Table 1 and Figure 2 summarize the results. Each non-Langevin baseline failed on at least one axis: bootstrap by construction ($\mathcal{N}{=}0$); Gaussian perturbation by negligible noise ($\sigma{\approx}0.003$, indistinguishable from bootstrap); and random convex combination by mode collapse ($\bar{\mathcal{D}}{=}0.008$). The DDPM baseline failed entirely: despite 5,000 training epochs, it produced samples with max-cosine similarity $0.062$ to the nearest stored pattern (indistinguishable from isotropic noise; cf. $0.057$ in the Gaussian control, Appendix L), confirming that diffusion models cannot learn the data distribution from K=100K{=}100 images. Among non-diffusion baselines, the VAE (latent dim 8, trained on the same K=100K{=}100 patterns) is the strongest: $\mathcal{N}{=}0.214\pm 0.005$, $\bar{\mathcal{D}}{=}0.441\pm 0.008$, surpassing GMM-PCA on both axes. At $\beta{=}200$ (SNR==0.036, the generation regime), SA dominates every baseline, including the best learned model (VAE), on both generation metrics simultaneously: novelty $0.548\pm 0.002$ ($2.6{\times}$ VAE, $2.8{\times}$ GMM-PCA), diversity $0.885\pm 0.002$ ($2.0{\times}$ VAE, $2.1{\times}$ GMM-PCA); energy rises above zero as samples explore off the attractor manifold (see ). To confirm that these high-novelty samples reflect structured generation rather than high-temperature noise, Appendix L compares SA at $\beta{=}200$ against Gaussian noise matched in per-pixel mean and variance: SA samples retain $45\%$ cosine similarity to the nearest stored pattern versus $33\%$ for the matched Gaussian and $6\%$ for isotropic noise, and have lower Hopfield energy ($1.47$ vs. $1.78$ and $2.34$), confirming that the sampler remains influenced by the memory geometry even in the generation regime. MALA’s 99.2% acceptance rate confirms negligible ULA bias at $\alpha{=}0.01$. A MALA comparison at $\beta{=}200$ (the generation regime) yields near-identical metrics: $\Delta\mathcal{N}{=}0.002$, $\Delta\bar{\mathcal{D}}{=}0.002$, $\Delta\bar{E}{=}0.018$ (acceptance $99.2\%$), confirming that the ULA bias is negligible across the full retrieval-to-generation spectrum (Appendix N). A step-size sweep (Appendix K) shows that ULA and MALA diverge only at $\alpha{\gtrsim}0.1$, and that MALA acceptance collapses to $0\%$ at $\alpha{\geq}0.2$ while ULA continues to produce samples with gracefully degrading quality. To test whether SA’s advantage over the VAE depends on the limited training data (K=100K{=}100), Appendix O compares against a VAE trained on $10{\times}$ more digit-3 images (N=1, [^0] N{=}1{,}000) and a VAE trained on all MNIST digits (N=10, [^0] N{=}10{,}000). At $\beta{=}2000$ (retrieval), SA achieves higher max-cosine similarity ($0.848$) than the best full-data VAE ($0.830$, all-MNIST, top-$150$ filtered) with $85\%$ higher diversity ($0.600$ vs. $0.325$), despite using $100{\times}$ less data and zero training. SA’s advantage is not an artifact of limited training data: it reflects the structural advantage of the exact, training-free score function. The SNR transition band (located via the entropy inflection of Proposition 3; $\mathrm{SNR}^{*}\approx 0.025$ for the d=64d{=}64 synthetic geometry, Section 4.3) thus delimits two qualitatively distinct regimes in the same algorithm: structured retrieval ($\beta{=}2000$, crisp digits near stored patterns) and genuine generation ($\beta{=}200$, blurry-but-recognizable novel digits). We emphasize that the 30-chain protocol at $\beta{=}2000$ is a structured retrieval strategy: each chain stays near its seed pattern, and the reported diversity of $0.600$ comes primarily from the multi-chain initialization, not from within-chain mixing (single-chain diversity at $\beta{=}2000$ is only $0.282$; Appendix F). The generation claim rests on $\beta{=}200$, where a single chain from a fixed seed achieves diversity $0.796$, exceeding the 30-chain $\beta{=}2000$ value, confirming genuine basin-crossing driven by the Langevin dynamics, not by initialization (Table 6, Appendix F).
To test whether DDPM’s failure is specific to the K=100K{=}100 regime, Appendix M reports a scaling study at $K\in\{100,\,500,\,1{,}000,\,2{,}000,\,3{,}500\}$ with $\beta$ set via entropy inflection at each $K$. DDPM’s max-cosine similarity never exceeds $0.09$ (near the isotropic noise floor), while SA achieves 0.130.130.180.18 in the generation regime and 0.240.240.330.33 in the retrieval regime. SA’s training-free score function provides a consistent structural advantage across the entire range tested.
Protein sequence generation.
To test SA on a non-image domain with discrete structure, we applied the full pipeline to protein sequences from the Pfam RRM family [24] (PF00076, RNA Recognition Motif): K=68K{=}68 aligned sequences, L=71L{=}71 positions, one-hot encoded ($\mathbb{R}^{1420}$), PCA-reduced to d=59d{=}59 (95% variance), and $\ell_{2}$-normalized. The entropy inflection (Proposition 3) yields $\beta^{*}{=}3.85$, half the random-pattern prediction $\sqrt{d}{=}7.68$, because conserved residues in the RRM motif increase the effective variance of query-key similarities. We ran SA at $\beta{=}77$ ($20\beta^{*}$, retrieval) and $\beta{=}8$ ($2\beta^{*}$, generation) with the same 30-chain protocol. Table 2 reports the key comparison against the VAE (full results in Appendix G).
Table 2: Protein sequence generation on Pfam PF00076 (K=68K{=}68, d=59d{=}59). KL: global amino acid composition divergence. Pos-KL: mean per-position KL divergence (captures conservation patterns at each residue). MI $r$: Pearson correlation of pairwise mutual information matrices (captures co-evolutionary coupling). HMM: fraction of generated sequences passing the Pfam RRM HMM filter at $E$-value <0.01<0.01. The generation regime ($\beta{=}8$) matches the VAE on novelty while achieving $3.4{\times}$ lower per-position divergence and $66\%$ higher coupling fidelity.
Method $\mathcal{N}$ $\uparrow$ SeqID KL $\downarrow$ Pos-KL $\downarrow$ MI $r$ $\uparrow$ HMM $\uparrow$
Bootstrap (replay) $0.000$ $0.644$ $0.143$ $7.52$ $0.692$ $100\%$
VAE (latent=8{=}8) $0.621$ $0.532$ $0.416$ $9.99$ $0.525$ $100\%$
SA ($\beta{=}77$, retrieval) $0.243$ $0.616$ $0.107$ $5.66$ $0.733$ $100\%$
SA ($\beta{=}8$, generation) $\mathbf{0.623}$ $0.538$ $\mathbf{0.060}$ $\mathbf{2.92}$ $\mathbf{0.871}$ $100\%$
At generation temperature, SA matches the VAE on novelty ($0.623$ vs. $0.621$) while achieving $6.9{\times}$ lower global amino acid composition divergence (KL=0.060=0.060 vs. $0.416$) and $3.4{\times}$ lower per-position KL divergence ($2.92$ vs. $9.99$). The per-position metric addresses a limitation of global composition: it computes the KL divergence at each aligned position independently and then averages, thereby capturing position-specific conservation patterns rather than just overall amino acid frequencies. SA-generated sequences also reproduce pairwise co-evolutionary coupling patterns with r=0.871r{=}0.871 Pearson correlation to the stored family’s mutual information matrix, compared to r=0.525r{=}0.525 for the VAE – a $66\%$ improvement. All generated sequences from every method pass the Pfam RRM hidden Markov model at $E$-value <0.01<0.01 (HMMER 3.4 [25]), with SA achieving median $E$-values of $8.1{\times}10^{-34}$ at generation temperature, confirming that the generated sequences are unambiguously recognized as RRM family members by an independent domain classifier. MALA at both temperatures ($\beta{=}77$ and $\beta{=}8$) produces near-identical results (acceptance rates $99.8\%$ and $99.8\%$; $\Delta\mathrm{KL}<0.002$), confirming negligible ULA bias on this domain. Full results including all baselines, MALA at generation temperature, and the per-position analysis are in Appendix G.
This is the paper’s clearest demonstration of the training-free score function’s structural advantage: because the Langevin gradient $\beta(\mathbf{T}(\boldsymbol{\xi})-\boldsymbol{\xi})$ is derived from the stored patterns at every step, it preserves family-level structure – at both the position-specific and pairwise coupling levels – even as samples explore far from individual stored sequences. The VAE, which must learn the distribution from K=68K{=}68 training examples, cannot maintain this fidelity when generating novel sequences. The temperature knob trades sequence identity ($0.616\to 0.538$) for novelty ($0.243\to 0.623$) while improving compositional fidelity (KL: $0.107\to 0.060$).
@@
Appendix E confirmed the same ranking on digits “1” and “8” (SA: $\mathcal{N}{\approx}0.152$$0.153$, $\bar{\mathcal{D}}{\approx}0.557$$0.591$; VAE: $\mathcal{N}{\approx}0.180$$0.209$, $\bar{\mathcal{D}}{\approx}0.397$$0.408$; MALA acceptance 99.2%): the VAE remains the strongest non-Langevin baseline on novelty but is far below both Langevin methods on diversity across all digit classes. Appendix H used S&P 500 log-returns (d=424d{=}424, K=2,766K{=}2{,}766) to characterize the novelty–fidelity trade-off on real data: SA achieved novelty $0.768\pm 0.001$ (bootstrap: $0.000$) while bootstrap led on marginal fidelity ($99.3\%$ vs. $64.2\%$ KS pass); volatility clustering was not reproduced, a principled consequence of fixed-$\beta$ equilibrium sampling. Appendix I scales to face images at d=4,096d{=}4{,}096: the SNR rule prescribes $\beta{=}10{,}000$; SA achieves novelty $0.159\pm 0.000$ and diversity $0.293\pm 0.001$, with ranking identical to MNIST across a $5.2{\times}$ dimension increase.
@@ 6 Discussion and Conclusion
The experiments validated the core claim: Langevin dynamics on the modern Hopfield energy converts deterministic attention into a principled stochastic sampler governed by a single temperature parameter. The temperature spectrum (Fig. 1a) exhibited a smooth phase transition between retrieval and generation, the continuous analogue of the classical Hopfield/Boltzmann duality [1] lifted to continuous states and Langevin dynamics. The convergence experiment (Fig. 1b) confirmed that Algorithm 1 reached the correct Boltzmann target, consistent with ULA theory [8]. The load-ratio sweep (Fig. 3) showed that generation quality degraded gracefully as $K/d$ increased, providing empirical evidence linking modern Hopfield capacity to generation quality.
The strongest quantitative result came from protein sequence generation (Table 2): on the Pfam RRM family (K=68K{=}68, d=59d{=}59), SA in the generation regime ($\beta{=}8$, $2\beta^{*}$) achieved $6.9{\times}$ lower global amino acid composition divergence than the VAE (KL=0.060{=}0.060 vs. $0.416$) and $3.4{\times}$ lower per-position KL divergence ($2.92$ vs. $9.99$) at matched novelty ($0.623$ vs. $0.621$). SA-generated sequences also reproduce pairwise co-evolutionary coupling patterns with r=0.871r{=}0.871 Pearson correlation to the stored family’s MI matrix, compared to r=0.525r{=}0.525 for the VAE. All generated sequences pass the Pfam RRM profile HMM at $E$-value <0.01<0.01 (median $8.1{\times}10^{-34}$), confirming biological validity. These results demonstrate the core advantage of the training-free score function at multiple levels of biological structure: because the Langevin gradient is derived from the stored patterns at every step, it preserves family-level fidelity in position-specific conservation and pairwise coupling, not just global composition. The entropy inflection condition correctly predicted a lower $\beta^{*}{=}3.85$ (half the random-pattern value $\sqrt{d}{=}7.68$), reflecting the increased effective similarity variance introduced by conserved residues in the RRM motif.
On MNIST digit images (Table 1), SA at generation temperature led the VAE by $2.6{\times}$ on novelty and $2.0{\times}$ on diversity, while matching the Metropolis-corrected MALA baseline ($99.2\%$ acceptance rate at $\alpha{=}0.01$). A DDPM baseline failed entirely at K=100K{=}100, producing samples indistinguishable from isotropic noise (max-$\cos{=}0.062$). A scaling study across $K\in\{100,\,500,\,1{,}000,\,2{,}000,\,3{,}500\}$ (Appendix M) confirmed that this failure persists across memory sizes: DDPM’s max-cosine never exceeds $0.09$, while SA maintains 0.130.130.180.18 in the generation regime and 0.240.240.330.33 in the retrieval regime. Appendices H and I extended these results to S&P 500 log-returns (d=424d{=}424) and face images (d=4,096d{=}4{,}096), confirming the SNR rule and the novelty–fidelity trade-off across a $5.2{\times}$ dimension range.
In summary, stochastic attention reuses the same primitive operations as a standard attention head and requires no learned score network, no training loop, and no contrastive objective. The analytic structure of the Hopfield energy (smoothness, dissipativity, and an exact closed-form gradient) yields a convergence guarantee in the convex regime ($\beta\sigma_{\max}^{2}<2$, Corollary 2) and geometric ergodicity of the continuous-time SDE for all $\beta>0$; at high $\beta$ the energy becomes multimodal and quantitative mixing bounds remain open, but the multi-chain protocol combined with fast intra-basin mixing provides a practical workaround (Section 4). The entropy inflection analysis (Proposition 3) provides a principled criterion for the retrieval–generation boundary: for random unit-norm memories the critical SNR scales as $d^{-1/4}$ (Eq. 14), recovering the empirically observed threshold of $0.025$ at d=64d{=}64. For structured data the effective variance of the similarities replaces $1/d$, but the inflection condition (12) remains a fully data-dependent selection rule for $\beta$.
@@
The ULA discretization introduces bias scaling with $\alpha$; our MALA comparison (Table 1) shows a 99.2% acceptance rate at $\alpha{=}0.01$, confirming negligible bias at this step size. At larger $\alpha$, MALA (Algorithm 2, Appendix C) is the preferred variant. At high temperature ($\beta{=}200$), SA samples are structured interpolations between stored patterns rather than crisp novel images: a Gaussian noise control (Appendix L) confirms that SA retains meaningful structural similarity to stored patterns (max-$\cos=0.453$ vs. $0.328$ for variance-matched noise), but sample quality is limited by the dominance of the noise term over the gradient. Higher-fidelity generation requires operating closer to the phase-transition band. As discussed in Section 4, inter-basin mixing time at high $\beta$ grows exponentially with the barrier height; quantitative scaling laws as a function of $\beta$ and $K/d$ remain to be established. Each step costs $\mathcal{O}(dK)$ (two matrix–vector products plus a softmax), identical to a single attention head; for d=784d{=}784, K=100K{=}100 the full 30-chain experiment completes in under 10 minutes on a laptop CPU. The method also requires $\mathbf{X}$ to be known and fixed at sampling time; streaming or latent memory settings would require extension. A toy multi-head extension using PCA-partitioned subspaces (Appendix Q) shows that the mechanism generalizes, but extension to learned multi-head projections, hierarchical memories, or full encoder-decoder transformers is future work.